Find the accumulate amount of

(a) $600 earns simple interest at 5% p.a. over 7 years.

(b) $480 compounded annually at 14% p.a. over 5 years

(c) $1,500 compounded semi-annually at 8% p.a. over 3 years.

(d) $2,000 compounded quarterly at 10% p.a. over 4 years.

(e) $500 compounded monthly at 12% p.a. over 6 years.


1
Expert's answer
2021-11-15T16:06:14-0500

a)

"I=PRT =600×\frac{5}{100}×7 = \$210"


b) compounded annually, n =1


"A=P(1+\frac{r}{n})^{nt}"


"=480(1+\frac{14}{100})^{5}= \$924.199"


c) compounded semi-annually,n=2


"A=P(1+\frac{r}{n})^{nt}"

"=1500(1+\frac{0.08}{2})^{2×3}"


"=1500(1.04)^6 = \$1897.979"


d) compounded quarterly,n=4


"A=P(1+\frac{r}{n})^{nt}"

"=2000(1+\frac{0.1}{4})^{4×4}"


"=2000(1.025)^{16} = \$2969.011"


e) compounded monthly, n=12


"A=P(1+\frac{r}{n})^{nt}"

"=500(1+\frac{0.12}{12})^{12×6}"


"=500(1.01)^{72}= \$1023.550"